Please help me solve this equation.

the variable expressions represent the angle measures of a triangle.

the measure of Angle A = x

The measure of angle B = 2x

The measure of angle c = 3x

Find the measure of each angle and classify the triangle by its angles.

Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

x+2x+3x=180
6x=180
x=30degrees
2x=60degrees
3x=90degrees
right triangle

Answer by stanbon(75887)

the variable expressions represent the angle measures of a triangle.
the measure of Angle A = x
The measure of angle B = 2x
The measure of angle c = 3x
Find the measure of each angle and classify the triangle by its angles.
-------------
Equation:
x + 2x + 3x = 6x = 180 degrees
x = 180/6 = 30 (A is 30 degrees (acute))
2x = 60 (B is 60 degrees (acute))
3x = 90 (C is 90 degrees (right angle))
=========================================
Cheers,
Stan H.